Details.

  • Telescope doen with the luci 2 mask exchange.
  • Dave confirma that we can start to use telescope.
  • Dave N30 camera alignment.
  • Power on 11:00am
  • ocam ok.
  • Memory issue. Clean up the disk, some core dumps were moved.
Deploy argos cal unit. not available.
  • checking status, not fullfil.
  • Reset hardware.
  • twice.
  • hardware available.

Check Bin3 rotation.

20220901_062538'

makePupils.movePupils('20200210_000000',0,-2,240,4) ->'20220901_063305'

makePupils.movePupils('20200210_000000',0,-3,240,4) ->20220901_070135

Set Flat

641 modes.

flat_20220905.png

Mon Sep 5 19:03:32.740 2022 aos.PresetAO{FLAO}.started - PresetAO {FLAO}[240] (aomode:ACE-AO,focst:bentGregorianFront,instr:LUCIFER,wfs:FLAOWFS,sox:0[0],soy:0,rox:0[0],roy:0,mag:14.7,cindex:0,,elev:nan,rotang:nan,grvang:nan,r0:0,skybr:nan,windsp:nan,winddir:nan)
Mon Sep 5 19:03:32.762 2022 aos.info - WFS source set to: FLAOWFS

test_062538 saturated at -100, error -0.9

test 063305

preset AOS again and it gaves an error.

Mon Sep 5 19:03:32.740 2022 aos.PresetAO{FLAO}.started - PresetAO {FLAO}[240] (aomode:ACE-AO,focst:bentGregorianFront,instr:LUCIFER,wfs:FLAOWFS,sox:0[0],soy:0,rox:0[0],roy:0,mag:14.7,cindex:0,,elev:nan,rotang:nan,grvang:nan,r0:0,skybr:nan,windsp:nan,winddir:nan)
Mon Sep 5 19:03:32.762 2022 aos.info - WFS source set to: FLAOWFS

close loop

Failure.

the first option the offloading was not running.

Open loop,

close loop and failure

Mon Sep 5 19:03:32.740 2022 aos.PresetAO{FLAO}.started - PresetAO {FLAO}[240] (aomode:ACE-AO,focst:bentGregorianFront,instr:LUCIFER,wfs:FLAOWFS,sox:0[0],soy:0,rox:0[0],roy:0,mag:14.7,cindex:0,,elev:nan,rotang:nan,grvang:nan,r0:0,skybr:nan,windsp:nan,winddir:nan)
Mon Sep 5 19:03:32.762 2022 aos.info - WFS source set to: FLAOWFS

adsc_stop/start

Anemometer information available.

preset 14.6

Mon Sep 5 19:03:32.740 2022 aos.PresetAO{FLAO}.started - PresetAO {FLAO}[240] (aomode:ACE-AO,focst:bentGregorianFront,instr:LUCIFER,wfs:FLAOWFS,sox:0[0],soy:0,rox:0[0],roy:0,mag:14.7,cindex:0,,elev:nan,rotang:nan,grvang:nan,r0:0,skybr:nan,windsp:nan,winddir:nan)
Mon Sep 5 19:03:32.762 2022 aos.info - WFS source set to: FLAOWFS

full close loop.

OK

b3_closeLoop_20220905.png

TN:_194756

wobble check.

12:50PM

check 070135

OK

bin3_20220905.png

ocst:bentGregorianFront,instr:LUCIFER,wfs:FLAOWFS,sox:0[0],soy:0,rox:0[0],roy:0,mag:14.6,cindex:0,,elev:nan,rotang:nan,grvang:nan,r0:0,skybr:nan,windsp:nan,winddir:nan)
Mon Sep 5 20:27:12.186 2022 aos.info - WFS source set to: FLAOWFS
  • load dist 1.5"
  • run automatic mode.
  • to be checked that the disturbance was injected. (?)
  • mag 15
  • The disturbance is not running......
TN:_203217

luci: Clear/H

20220906.00019,20

modal SR 60

bin3_psf_20220905.png

Bin4

Mon Sep 5 20:39:23.556 2022 aos.PresetAO{FLAO}.started - PresetAO {FLAO}[255] (aomode:ACE-AO,focst:bentGregorianFront,instr:LUCIFER,wfs:FLAOWFS,sox:0[0],soy:0,rox:0[0],roy:0,mag:16.6,cindex:0,,elev:nan,rotang:nan,grvang:nan,r0:0,skybr:nan,windsp:nan,winddir:nan)
Mon Sep 5 20:39:23.571 2022 aos.info - WFS source set to: FLAOWFS

144PM check

pupil:20220831_061714

tn:205731

pupil:20200916_173857

better flat pupil.

tn:_210224

b4_pupil20200916_173857.png

after loading the pupil 20220831_061714 again better flat.

TN:_210624

b4-pupil20220831_061714.png

loaded ao table with pupil B4: 20220831_061714

bin1

Mon Sep 5 21:20:45.494 2022 aos.PresetAO{FLAO}.complete - PresetAO {FLAO}[262] - dx:nan dy:nan slNull: f1:0.0 f2:1.0 freq:452 mag:0 nBins:4 nModes:54 r0:nan snMode:0 strehl:0 ttMod:0
Mon Sep 5 21:21:01.389 2022 aos.PresetAO{FLAO}.started - PresetAO {FLAO}[263] (aomode:ACE-AO,focst:bentGregorianFront,instr:LUCIFER,wfs:FLAOWFS,sox:0[0],soy:0,rox:0[0],roy:0,mag:10,cindex:0,,elev:nan,rotang:nan,grvang:nan,r0:0,skybr:nan,windsp:nan,winddir:nan)
Mon Sep 5 21:21:01.410 2022 aos.info - WFS source set to: FLAOWFS

loadedloaded dist 1.5"

it doen;t appear in the AOS log

b1

it seems some actuators are jumping during the close loop. To analysie the telemetry

bin12022-09-05 at 2.27.23.png

_212741,_212847

The PSf shows to be oscillating. Analysis of the TN.

PSF_b1-202209005_212847.png

final hexapod.

2.33,1.29,-1.38,137,-187,0

Memory.

[AOeng@soul-sxwfs source]$ df -ah

Filesystem Size Used Avail Use% Mounted on

/dev/mapper/vg00-root

5.9G 5.4G 255M 96% /

proc 0 0 0 - /proc

sysfs 0 0 0 - /sys

devpts 0 0 0 - /dev/pts

tmpfs 7.8G 12K 7.8G 1% /dev/shm

/dev/sda1 190M 40M 141M 22% /boot

/dev/mapper/vg00-home

788G 657G 92G 88% /home

/dev/mapper/vg00-tmp 976M 12M 914M 2% /tmp

/dev/mapper/vg00-usr 5.8G 5.1G 487M 92% /usr

/dev/mapper/vg00-var 3.9G 702M 3.0G 19% /var

/dev/mapper/vg00-opt 488M 241M 223M 52% /opt

none 0 0 0 - /proc/sys/fs/binfmt_misc

sunrpc 0 0 0 - /var/lib/nfs/rpc_pipefs

192.168.39.20:/volume3/lbto_flao

        1. T 16T 2.0T 90% /ao-data

192.168.39.54:/local/aomeas

7.2T 5.7T 1.2T 84% /local/aomeas

archive-in:/mnt/newdata

        1. T 1.2T 9.6T 11% /mnt/newdata

[AOeng@soul-sxwfs source]$

AO tables for ACE and TTM modified for pupils Bin3 and Bin4.

AO table loaded:
  • table_LUCIFER_TTM-AO.txt.20220902
  • table_LUCIFER_ACE-AO.txt.20220902

2:28 PM stop

AO data.

20220905_212741

high atteniation of SR; due to TT jitter 0.59502592

tt_atteniation_20220905_212741.png

-- JuanCarlosGuerra - 05 Sep 2022
Topic revision: r4 - 13 Sep 2022, JuanCarlosGuerra
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback